OpenStudy (anonymous):

Rewriting Expressions.... Such as: 1. 5 more than x 2. The sum of two numbers is 12, if one number is x what is the other number.

OpenStudy (anonymous):

y=5+x x+y=12 x+5+x=12 x=7/2 y=17/2

OpenStudy (anonymous):

the other is y

OpenStudy (anonymous):

Thanks!